OpenStudy (anonymous):

How many grams of Sodium carbonate (Na2CO3) contain 1.773x10^17 carbon atoms?

OpenStudy (anonymous):

HOw do i solve this?

OpenStudy (anonymous):

1 mole of sodium carbonate = 106 g ..As sodium carbonate contain 1 carbon atom = 6.022 x 10 ^ 23 so 1.773 x 10 ^17/ 6.022 x 10 ^23 = ur answer

OpenStudy (anonymous):

Na2CO3 m= ? N =1.773 x 10^17 First, you need to find moles using the N (number of carbon atoms) and Avogadro's number (6.022 x 10^23) and this formula: n= N / NA Using the moles, you can figure out mass by multiplying the moles by the molar mass of the compound. Ta da.